NOAA Office of Satellite and Product Operations

NOAA 16 AVHRR Subsystem Summary

Component Description Status
CHANNEL 1Spectral channel #1, 0.58-0.68 micro meters. Green
CHANNEL 2Spectral channel #2, 0.725-1.0 micro meters. Green
CHANNEL 3ASpectral channel #3A, 1.58-1.64 micro meters. Green
CHANNEL 3BSpectral channel #3B, 3.55-3.93 micro meters. Green
CHANNEL 4Spectral channel #4, 10.3-11.3 micro meters. Green
CHANNEL 5Spectral channel #5, 11.5-12.5 micro meters. Green
OPTICSAVHRR Optical mirrors and telescopesGreen
PSUPower Supply UnitGreen

Subsystem History:

Date Time Component Description
10/18/200516:30:00SCAN MOTORAVHRR scan motor current significantly improved over past several days as is synch delta values and imagery. No bar coded images since 12 Oct. Status upgraded to YELLOW.
10/06/200503:00:00SCAN MOTORAVHRR scan motor current spiking and causing intermittent imagery bar codes again. Imagery clears up with each MIRP Rephase enable operation, executed once per orbit over the south pole but frequently returns a few minutes later. This is a chronic problem with this payload.
10/03/200514:00:00SCAN MOTORN16 AVHRR scan motor current significantly improved as are synch delta values. Imagery bar codes ceased.
09/29/200516:00:00SCAN MOTORN16 AVHRR downgraded to RED status because of elevated scan motor currents and persistent bar coded imagery.
09/24/200518:00:00SCAN MOTORAVHRR scan motor current higher and less stable. Imagery bar codes starting to intermittently appear. Imagery clearing up with each MIRP Rephase enable operation, executed once per orbit over the south pole.
09/01/200519:30:00SCAN MOTORAVHRR Bar code imagery started on 31 Aug at 01:15z and persisted through a manual" MIRP rephase enable command at WCDA on 1 Sept at 19:30Z. Cause of bar code persistence: MIRP rephase enable commands (usually executed out of stored commands once per orbit over the South Pole) were omitted from stored command table starting 30 Aug. To be reinserted 2 Sept."
06/29/200507:00:00SCAN MOTORAVHRR bar coded imagery occurred for the first time since 22 May 05 in conjunction with scan motor current exceeding 280 mAmps. Cleared up after MIRP rephase over South Pole approximately 25 minutes later and a reduction in scan motor current belwo 280.
05/22/200517:45:00SCAN MOTORAVHRR scan motor surged to 310 mAmps causing bar coded imagery for the first time since early August 2004. Imagery cleared after MIRP rephase occured over south pole on next orbit.
10/23/200412:00:00SCAN MOTORAVHRR scan motor current became elevated. Synch deltas also quickly runup to pegged" after each rephase over south pole. Imagery degraded (waves) but no bar codes as yet."
10/01/2004SCAN MOTORAVHRR Scan motor current improved back to pre-anomalous values, synch deltas also nominal. Imagery clear
06/04/200420:00:00SCAN MOTORAVHRR scan motor current, synch delta and imagery continues to have wide fluctuations in quality. Highly variable and unpredictable.
05/19/200412:01:00 Enable MIRP rephasing for 10 seconds out of stored commands each orbit over the south pole to reduce percentage of bar coded imagery.
04/27/200414:00:00 Begin implementing intermittent manual" MIRP rephasing when AVHRR bar coded imagery is apparent to clean up the images. "
03/20/200418:50:00SCAN MOTORScan motor intermittent current surges similar to 16Mar04 profile resumed. Most images showing bar codes.
03/18/200403:40:00SCAN MOTORAVHRR Scan motor broke through an apparent obstruction. Quick surge in motor current and temp, followed by a slow decline to pre-incident values. Imagery barcodes cleared up after MIRP rephased at 04:30Z
03/16/200402:46:00SCAN MOTORAVHRR Scan motor current began intermittently surging to 300+ mAmps. Imagery barcodes started to occur. Surges lasted approx 2 minutes then retreated back to 265 mA for 15 minutes then surged again, continually.
01/24/2004SCAN MOTORAVHRR HRPT images no longer showing barcode patterns. Scan motor current slowly improving (trending lower).
01/16/200412:20:00SCAN MOTORTCE24 (AVHRR heater and louver) commanded back on. AVHRR component temps dropped an average of 4 degs C within 2 hours.
01/15/200419:02:00ELECTRONICSAVHRR/MIRP rephase DISABLE.
01/15/200419:01:00ELECTRONICSAVHRR/MIRP rephase enable.
01/14/200413:55:00SCAN MOTORAVHRR Scan motor current surged to maximum reported value (305 mAmps). Imagery began showing bar code" patterns shortly thereafter."
01/06/2004SCAN MOTORAVHRR declared YELLOW in status because of degraded imagery, scan motor performance and synch delta statistics.
01/06/2004 AVHRR Imagery showing wavey patterns at image edges.
01/04/2004SCAN MOTORAVHRR Scan Motor and synch delta at significantly elevated levels, imagery exhibiting wave patterns.
10/17/2003SCAN MOTORAVHRR scan motor current continuously nominal for over 2 weeks. Upgrade status to Green
09/29/200302:15:00SCAN MOTORAVHRR scan motor current, synch delta values and imagery returned to nominal Maintaining yellow status for the time being.
09/26/200304:30:10SCAN MOTORAVHRR / MIRP Rephase disable
09/26/200304:30:00SCAN MOTORAVHRR / MIRP Rephase enable
09/23/200317:05:00SCAN MOTORAVHRR rephase enable.
09/23/200317:05:00SCAN MOTORAVHRR rephase disable
09/23/200314:00:00SCAN MOTORTurned off TCE 24 (AVHRR heater and Louver) to try and warm AVHRR and reduce scan motor lubricant viscosity.
09/18/200318:00:00SCAN MOTORAVHRR Rephase Disable because of scan motor anomaly. Rephase 2x daily.
09/17/200316:40:00SCAN MOTORAVHRR scan motor current and synch delta values acting erratically. Status downgraded to YELLOW.
08/01/200322:10:00 Activates Macro F, ascending AVHRR channel 3A macro
07/28/200313:00:00 Activates Macro F, ascending AVHRR channel 3A macro
07/13/200312:16:00CHANNEL 3AActivation of macro F for polar region AVHRR channel 3A use. 3A active between 65 N (magnetic) and peak of orbit.
05/01/200316:05:00OPTICSSwitch to channel 3B only, inhibit automatic 3A-B switching.
01/16/200312:24:00ELECTRONICSDisable MIRP/AVHRR rephaseing
01/16/200312:23:00ELECTRONICSEnable MIRP/AVHRR rephaseing
07/03/200212:01:00CHANNEL 3BDisable AVHRR Channel 3B macro
06/15/200218:00:00CHANNEL 3BUpload and enable macro for switching to AVHRR Channel 3B over Colorado fires
07/26/2001 1 second ground system time measurement software error found - caused along-track error
06/01/2001 Investigation determined instrument alignment not cause of along-track error
09/22/2000 AVHRR 1 second along-track pointing error observed
09/21/200010:56:00PSUOn Orbit
09/21/200010:56:00SCAN MOTOROn Orbit
09/21/200010:56:00ELECTRONICSOn Orbit
09/21/200010:56:00RADIANT COOLEROn Orbit
09/21/200010:56:00OPTICSOn Orbit
09/21/200010:56:00BASEPLATEOn Orbit
09/21/200010:56:00CHANNEL 1On Orbit
09/21/200010:56:00CHANNEL 2On Orbit
09/21/200010:56:00CHANNEL 3AOn Orbit
09/21/200010:56:00CHANNEL 3BOn Orbit
09/21/200010:56:00CHANNEL 4On Orbit
09/21/200010:56:00CHANNEL 5On Orbit